# A Superconducting Single-Atom Phonon Laser

### v2

@@
A Superconducting Single-Atom Phonon Laser
[…]
@@ A Superconducting Single-Atom Phonon Laser
The development of quantum acoustics has enabled the cooling of mechanical objects to their quantum ground state, generation of mechanical Fock-states, and Schrödinger cat states. Such demonstrations have made mechanical resonators attractive candidates for quantum information processing, metrology, and macroscopic tests of quantum mechanics. Here, we experimentally demonstrate a direct quantum-acoustic equivalent of a single-atom laser. A single superconducting qubit coupled to a high-overtone bulk acoustic resonator is used to drive the onset of phonon lasing. We observe the absence of a sharp lower lasing threshold and characteristic upper lasing threshold, unique predictions of single-atom lasing. Lasing of an object with a $25$ $\mu$g mass represents a new regime of laser physics. It provides a possible tool for generating large amplitude coherent states in circuit quantum acoustodynamics, which is important for fundamental and quantum information applications.
@@ INTRODUCTION
The invention of the laser in the early 1960s has proven to be one of the most important breakthroughs in modern physics. The laser found its niche in the advancement of atomic physics, resulting in numerous groundbreaking discoveries such as spectroscopic measurements of atomic species Schawlow (1982, 1978), laser cooling of atoms Phillips and Metcalf (1982); Phillips (1998), and the realization of Bose-Einstein condensation Anderson et al. (1995); Davis et al. (1995). Beyond this, lasers are indispensable tools in areas like gravitational wave detection Abbott et al. (2016), biology Chaurasiya et al. (2016), chemistry McCreery (2005), and even in medicine Khalkhal et al. (2019). The advancements enabled by optical lasers have motivated the development of lasing states of mechanical vibrations — phonon lasersMahboob et al. (2013); Vahala et al. (2009); Pettit et al. (2019); Kuang et al. (2023); Behrle et al. (2023). Recent examples of phonon lasing have been realized with the center of mass motion of levitated microspheres or ions Vahala et al. (2009); Pettit et al. (2019); Kuang et al. (2023); Behrle et al. (2023), limiting the integrability of these coherent phonon sources. In other examples of phonon lasers in literature, the lasing state is the result of dispersive or dissipative optomechanical coupling Shandilya et al. (2021); Cui et al. (2021); Zhang et al. (2018); Grudinin et al. (2010); Potts et al. (2021); Parsa et al. (2024) which are well described as mechanical parametric instabilities Kippenberg and Vahala (2008) or via electron transport driven Kent et al. (2006) and optically driven semiconductor heterostructures Bron and Grill (1978); Wallentowitz et al. (1996); Camps et al. (2001); Tilstra, Arts, and De Wijn (2007).
The prospect of controlling and manipulating phonons at the quantum level has sparked the development of the field of circuit quantum acoustodynamics (cQAD), in which single superconducting qubits are coupled to bulk acoustic phonon modes Chu et al. (2017). Circuit QAD has seen rapid success with demonstrations such as the generation of arbitrary Fock states Chu et al. (2018), and Schrödinger cat states Bild et al. (2023), and the prospect of developing novel hybrid quantum systems Chu and Gröblacher (2020); Välimaa et al. (2022); Hann et al. (2019). However, as the acoustic mode is coupled only via the qubit in current cQAD platforms, it can be challenging to generate the large coherent phonon amplitudes required for creating large cat states and for phonon mode tomography Bild et al. (2023).
Figure 1: Schematic of the on-chip $\hbar$BAR device. Rendering of the $\hbar$BAR device. The $\hbar$BAR device comprises two chips bounded in a flip-chip orientation. The top chip is 650 $\mu$m of sapphire, hosts the HBAR modes (red), and is coupled to the superconducting antenna using an aluminum nitride pad (red). The pocket-style transmon qubit (silver) is fabricated from NbTiN on the bottom silicon chip and coupled to the feedline via an on-chip readout resonator (blue).
Figure 2: Qubit-induced acoustic linewidth narrowing. a Definition of modes and the coupling rates between the modes. The readout resonator $\hat{a}$ has input-output modes labelled $1$ and $2$. b Measured two-tone spectroscopy for drive power of -12.0 dBm set at room temperature. c Measured phonon-induced transparency window as a function of qubit drive power. Starting from the lowest curve, drive powers are -12.0, -6.0, and -2.0 dBm set at room temperature, respectively. With increasing power, two features can be noticed. The transparency window reduces in depth, and the full-width half-maximum linewidth narrows. d Extracted experimental full-width half-maximum of the transparency window as a function of qubit drive power. Colored points match the corresponding curves in c.
@@
In this article, we demonstrate the generation of large coherent phonon populations in cQAD using a direct phononic analog of a single-atom optical laser McKeever et al. (2003); Dubin et al. (2010); An et al. (1994). A single two-level artificial atom couples to a bulk on-chip phonon mode, generating a large amplitude phonon coherent state through the process of single-atom lasing. Similar to conventional lasers, in single-atom lasers, the excitations in the lasing mode are dominantly produced by stimulated emission Mu and Savage (1992), enabling a potential route to large coherent state preparation in cQAD. The lasing is observed through a dramatically extended ringdown of the artificial atom excited by the phonon laser and demonstrates the nearly threshold-free nature of single-atom lasers. We also observe a unique predicted feature of single-atom lasers, an upper lasing threshold Dubin et al. (2010); Mu and Savage (1992); Ashhab et al. (2009). Our experimental observations agree with the theoretical and numerical models we use to characterize the system.
@@
Finally, the phonon mode’s bulk nature may allow the integration of color centers or quantum dots, enabling strain engineering or coherent mechanical driving for future hybrid quantum systems. Moreover, the coherent state generated via phonon lasing may be used as an efficient displacement pulse for generating large Schrödinger cat states Bild et al. (2023) or squeezed states Marti et al. (2024). Generating large mechanical cat states has exciting potential applications for future macroscopic tests of quantum mechanics Schrinski et al. (2023).
@@
The qubit’s state was measured by applying a weak probe tone on resonance with the readout resonator and monitoring the transmitted signal as a second tone was swept near the qubit frequency. The transmission spectrum at the readout frequency directly maps to the qubit occupation $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$; see Fig. 2(b). The asymmetry of the qubit spectrum is due to the finite photon population within the readout resonator Gambetta et al. (2006); Lachance-Quirion et al. (2017) and is well described by our theoretical model. Moreover, the narrow transparency window within the qubit spectrum results from the weak hybridization between the qubit and the phonon mode. In this work, the qubit and the phonon frequency were detuned by approximately $3.3$ MHz. See the discussion in the supplementary text for full details. However, as will be discussed below, coherent energy exchange between the qubit and phonon modes do not provide a clear picture of the dynamics at high drive powers.
PHONON LASING
Three distinct features can be observed within the two-tone spectrum when increasing the qubit drive power. First, the qubit linewidth is power-broadened Schuster et al. (2005); Gambetta et al. (2006). Large qubit drive powers increase the qubit decay rate due to increased stimulated emission. Less intuitive is the gradual disappearance and narrowing of the phonon-induced transparency window; see Fig. 2(c). The total linewidth of the transparency window is proportional to the phonon linewidth and demonstrates an inverse dependence on the drive power; see Fig. 2(d). This dependence is reminiscent of linewidth narrowing described by the Schawlow-Townes equation and is a known property of lasing states Gordon, Zeiger, and Townes (1955). However, it is important to note that, here, we observe a narrowing of the induced transparency window and not a narrowing of the emission spectrum. These features can be understood by considering the schematic shown in Fig. 3(a).
The build-up of the phonon population can be understood by considering the different processes through which excitations can be transferred between the phonon and the qubit. At low driving powers, the linewidth broadening of the qubit due to the microwave drive is small, and the phonon mode and qubit hybridize weakly. The weak hybridization combined with the microwave drive initiates the building up of the phonon population. A balance between the rate at which the qubit is excited and de-excited, the rate at which excitations are transferred between qubit and phonon mode and the stochastic phonon losses yield a steady-state occupation of the phonon mode. For low drive powers and thus slow qubit excitation rates, phonon loss dominates, preventing the build-up of the phonon population. As the qubit excitation rate increases, transitions between states of the Jaynes-Cummings ladder with higher phonon numbers can be efficiently driven. Since the rate at which excitations are transferred between states of the Jaynes-Cummings ladder depends on the phonon number of the involved states, as the average population of the phonon mode increases, the rate at which excitations are transferred correspondingly increases; this is the origin of acoustic stimulated emission Hofheinz et al. (2008). In the steady state, this process, balanced by stochastic phonon losses, results in a phonon mode described by a large coherent state, i.e. the phonon mode starts lasing. Such processes also provide an intuitive understanding of the reduced visibility of the transparency window. As the drive power increases, the linewidth of the qubit becomes increasingly broad, and the coupling between qubit and phonon is not strong enough to overcome the qubit loss; therefore, the qubit spectroscopy provides limited information about the phonon mode.
@@
Figure 3: Phonon lasing observed through qubit ringdown dynamics. a Schematic of the energy levels for the weakly hybridized phonon-qubit system. The qubit acts as an artificial two-level atom, with ground state $|g\rangle$ and excited state $|e\rangle$. A coherent drive of strength $\Omega_{\rm q}$ drives the qubit between its ground and excited state. The rate at which excitation transfer processes $|e,N-1\rangle\rightarrow|g,N\rangle$ occur scales as $\sqrt{N}$, where $N$ is the total number of excitations, while the phonon relaxation scales linearly with $N$. The qubit is rapidly re-excited for strong pump powers, resulting in a build-up of phonon excitations. This process results in phonon lasing. b Measured qubit ringdown for a qubit drive of -3.0 dBm. Blue line: the qubit drive is detuned from the HBAR by $250$ kHz with the phonon mode in the non-lasing state, decaying on a time scale of the $\sim 200$ ns decay of the qubit. Red line: The qubit drive is tuned directly on resonance with the HBAR mode, exciting it into the lasing state, exhibiting a dramatically longer, non-exponential decay due to re-excitation from the coherently excited phonon mode. The master equation simulation (dashed black line) is plotted over the data.
Figure 4: Unique signatures of single-atom lasing. a Measured gated qubit ringdown for a qubit drive of -12.0, -8.0, -6.0, and -2.0 dBm. With increasing drive power, the gated qubit ringdowns increase in amplitude and duration as the phonon mode population increases. b Measured gated qubit ringdown for a qubit drive of -2.0, 2.0, 4.0, and 8.0 dBm. The qubit drive is tuned directly on resonance with the HBAR mode for all measurements. With increasing drive power, the gated qubit ringdowns decrease in amplitude and duration as the phonon mode population decreases above the self-quenching threshold. For all measurements in a and b, the qubit drive is tuned directly on resonance with the HBAR mode. c Simulated phonon state population $\langle\hat{b}^{\dagger}\hat{b}\rangle$ as a function of qubit drive power. The phonon population is at its maximum at a drive power of approximately -2.0 dBm and decreases for drive powers above the upper lasing threshold. The colored data points indicate the corresponding trace colors in a and b.
@@
At the highest qubit drive powers, the qubit linewidth has been increased such that the rapid decay of the qubit results in a self-quenching of phonon lasing Mu and Savage (1992). This results in an upper lasing threshold above which the phonon mode is no longer lasing. Above this threshold, the qubit undergoes rapid Rabi oscillations due to the strong microwave drive and, therefore, cannot exchange excitations with the phonon mode, reducing the phonon amplitude. It should also be noted that the phonon mode statistics are no longer described as a coherent state above the upper threshold. This process has been described previously in the context of single-atom photon lasers McKeever et al. (2003) and is captured by our theoretical description; see the supplementary text.
@@
Directly probing the phonon mode is not possible in the current experimental configuration since the readout is performed via the two-level system and not through a propagating photon mode Kuang et al. (2023); Behrle et al. (2023). Direct measurements of the Rabi oscillations between the qubit and phonon state have been previously used to measure Fock-states in $\hbar$BAR devices Chu et al. (2018). However, this data would not be possible due to the short lifetime of our qubit and the multiplicity of simultaneous Rabi oscillation frequencies between two Fock states is given by $2g_{\rm N}=2\sqrt{N}g_{\rm qb}$, which scale with the phonon Fock number $N$Chu et al. (2018). Instead, we rely on the mismatch between the decay rate of the phonon mode $\tau\sim 25$ $\mu$s and that of the qubit. Using gated two-tone ringdown measurements, we can distinguish pure qubit decay from qubit decay driven by a highly excited phonon state. If the phonon mode is highly excited—in the absence of an external drive—the coherent Jaynes-Cummings interaction will continually drive the qubit, resulting in an extended relaxation of the qubit population compared to its intrinsic relaxation rate.
@@
First, the drive power was set near the peak of the lasing amplitude and was detuned $250$ kHz above the HBAR resonance. The gated two-tone measurement was performed, and the blue data points in Fig. 3(b) show the resulting time domain measurement and a ringdown on the order of $\sim 200$ ns. This corresponds to the intrinsic $T_{1}$ decay of the transmon qubit. A second measurement was performed at the same drive power, but the drive was tuned resonant with the HBAR. A ringdown on the order of $25$ $\mu$s is observed for this configuration, represented by the red data points in Fig. 3(b). The extended ringdown confirms the highly excited phonon amplitude of the mechanical state. When the qubit drive is switched off, the phonon mode can exchange excitations with the qubit, continually re-exciting the qubit until the phonon mode has decayed back to its ground state. Moreover, near the peak phonon amplitude, in contrast to a thermal state, the coherent state generated by the lasing results in a qubit ringdown that is not exponential; instead, the qubit ringdown is approximately linear. This feature is captured by our numerical model, the dashed curve in Fig. 3(b). Moreover, our model also captures the ring-up of the qubit and is described within the supplementary information.
@@
We can estimate the phonon population from our numerical model by fitting the spectroscopic and ringdown data. The estimated phonon population is shown in Fig. 4(c) as a function of qubit drive power. At a power of -2.0 dBm, the phonon population is nearly maximized, corresponding to the ringdown in Fig. 4(a,b), and the upper lasing threshold is clearly visible as the phonon population rapidly reduces with increasing drive power. The upper threshold is experimentally confirmed by performing a set of gated ringdown measurements at a series of qubit drive powers. With increasing qubit drive power, the individual ringdown traces grow in amplitude and increase in duration, corresponding to the increasing phonon population, as shown in Fig. 4(a). At a drive power of -2.0 dBm, the phonon ringdown obtains its peak amplitude and duration, indicating a peak in the phonon population, shown in Fig. 4, which agrees with the minimum in the transparency window linewidth, shown in Fig. 2(b). Further increasing the qubit drive power beyond -2.0 dBm the qubit ringdown decreases in amplitude and duration, a direct indication of the self-quenching nature of the single-atom phonon laser, see Fig. 4(b). The upper threshold is a defining signature of a single-atom laser Mu and Savage (1992), and the agreement between the gated qubit ringdown and our numerical simulations indicates that we have indeed demonstrated a single-atom phonon laser.
CONCLUSION

### v3

# Generation of Large Amplitude Phonon States in Quantum Acoustics

@@
Generation of Large Amplitude Phonon States in Quantum Acoustics
Clinton A. Potts1†, Wilfred J.M. Franse1, Victor Augusto S.V. Bittencourt2, Anja Metelmann2,3,4 and Gary A. Steele1∗
@@
2Institut de Science et d’Ingénierie Supramoléculaires (ISIS, UMR7006), Université de Strasbourg and CNRS, 67000 Strasbourg, France
@@
The development of quantum acoustics has enabled the cooling of mechanical objects to their quantum ground state, generation of mechanical Fock-states, and Schrödinger cat states. Such demonstrations have made mechanical resonators attractive candidates for quantum information processing, metrology, and macroscopic tests of quantum mechanics. However, generating large-amplitude phonon states in quantum acoustic systems has been elusive. In this work, a single superconducting qubit coupled to a high-overtone bulk acoustic resonator is used to generate a large phonon population in an acoustic mode of a high-overtone resonator. We observe extended ringdowns of the qubit, confirming the generation of a large amplitude phonon state, and also observe an upper threshold behavior, a consequence of phonon quenching predicted by our model. This work provides a key tool for generating arbitrary phonon states in circuit quantum acoustodynamics, which is important for fundamental and quantum information applications.
@@ Introduction
The prospect of controlling and manipulating phonons at the quantum level has sparked considerable interest, in particular for applications of non-classical phonon states for quantum sensing Fadel (2023); Pino et al. (2018); Zivari et al. (2022). The latter might enable tests of gravity effects in quantum mechanics, a long-pursued milestone in physics Penrose (1996); Diosi (1987); Gely and Steele (2021). In this context, circuit quantum acoustodynamics (cQAD)Chu et al. (2017) has emerged as a promising platform for controlling phonons at the quantum regime. In such systems, a single superconducting qubit is coupled to an acoustic phonon mode, enabling the use of the well-developed circuit quantum electrodynamics toolkit Blais et al. (2021); Krantz et al. (2019); Hofheinz et al. (2009) for manipulating qubits as a way to control phonons Undershute et al. (2025); Franse et al. (2024); Manenti et al. (2017); Jiang et al. (2023). As a consequence, cQAD has seen rapid success in the generation of mechanical states including the generation of Fock states Chu et al. (2018) and Schrödinger cat states Bild et al. (2023). Moreover, cQAD has prospects for developing novel hybrid quantum systems Chu and Gröblacher (2020); Välimaa et al. (2022); Hann et al. (2019), allowing further integration of the system with other hybrid architectures Clerk et al. (2020).
@@
In this article, we demonstrate the generation of a large phonon population in cQAD using a scheme similar to single-atom lasers McKeever et al. (2003); Dubin et al. (2010); An et al. (1994). A single superconducting qubit couples to a bulk on-chip phonon mode, generating a large amplitude phonon state through the process of stimulated emission. The phonon state amplitude is confirmed through a dramatically extended ringdown of the superconducting qubit excited by the phonon state. We also observe a unique predicted feature of our model, an upper threshold, which is a consequence of population quenching, a phenomenon also present in single-atom lasers Dubin et al. (2010); Mu and Savage (1992); Ashhab et al. (2009). Close to such an upper threshold, the phonon mode experiences a reduced linewidth, indicating a qubit-induced amplification. Our experimental observations agree with the theoretical and numerical models we use to characterize the system.
Figure 1: Schematic of the on-chip $\hbar$BAR device. Rendering of the $\hbar$BAR device. The $\hbar$BAR device comprises two chips bonded in a flip-chip orientation. The top chip is 650 $\mu$m of sapphire, hosts the high-overtone bulk acoustic wave resonances (HBAR) modes (red), and is coupled to the superconducting antenna using an aluminum nitride pad (red). The pocket-style transmon qubit (silver) is fabricated from niobium titanium nitride on the bottom silicon chip and coupled to the feedline via an on-chip readout resonator (blue). Art produced by Enrique Sahagun Sci .
@@
Finally, the phonon mode’s bulk nature may allow the integration of color centers or quantum dots, enabling strain engineering or coherent mechanical driving for future hybrid quantum systems. Moreover, the coherent state generated here may be used as an efficient displacement pulse for generating large Schrödinger cat states Bild et al. (2023) or squeezed states Marti et al. (2024). Generating large mechanical cat states has exciting potential applications for future macroscopic tests of quantum mechanics Schrinski et al. (2023).
Results
@@
The qubit’s state was measured by applying a weak probe tone on resonance with the readout resonator and monitoring the transmitted signal as a second tone was swept near the qubit frequency. The transmission spectrum at the readout frequency directly maps to the qubit occupation $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$; see Fig. 2(b). The asymmetry of the qubit spectrum is due to the finite photon population within the readout resonator Gambetta et al. (2006); Lachance-Quirion et al. (2017) and is well described by our theoretical model. Moreover, the narrow transparency window within the qubit spectrum results from the weak hybridization between the qubit and the phonon mode. In this work, the qubit and the phonon frequency were detuned by approximately $3.3$ MHz. See the discussion in the supplementary text for full details. However, as will be discussed below, coherent energy exchange between the qubit and phonon modes does not provide a clear picture of the dynamics at high drive powers.
Figure 2: Qubit-induced acoustic linewidth narrowing. a Definition of modes and the coupling rates between the modes. The readout resonator $\hat{a}$ has input-output modes labelled $1$ and $2$, and is dispersively coupled to the qubit $\hat{\sigma}_{\rm z}$ with a rate $\chi$. The mechanical mode is labeled $\hat{b}$ and has a qubit-phonon coupling rate $g_{\rm qb}$. b Measured two-tone spectroscopy for drive power of -12.0 dBm set at room temperature. c Measured phonon-induced transparency window as a function of qubit drive power. Starting from the lowest curve, drive powers are -12.0, -6.0, and -2.0 dBm set at room temperature, respectively. With increasing power, two features can be noticed. The transparency window reduces in depth, and the full-width half-maximum linewidth narrows. d Extracted experimental full-width half-maximum of the transparency window as a function of qubit drive power. Colored points match the corresponding curves in c.
@@
Figure 3: Large amplitude phonon states observed through qubit ringdown dynamics. a Schematic of the energy levels for the weakly hybridized phonon-qubit system. The qubit acts as an artificial two-level atom, with ground state $|g\rangle$ and excited state $|e\rangle$. A coherent drive of strength $\Omega_{\rm q}$ drives the qubit between its ground and excited state. The qubit-mechanical couple at a rate $g_{\rm qb}$ and a qubit and mechanics decay at a rate $\gamma_{\rm q}$ and $\gamma_{\rm b}$, respectively. The rate at which excitation transfer processes $|e,N-1\rangle\rightarrow|g,N\rangle$ occur scales as $\sqrt{N}$, where $N$ is the total number of excitations, while the phonon relaxation scales linearly with $N$. The qubit is rapidly re-excited for strong pump powers, resulting in a build-up of phonon excitations. b Measured qubit ringdown for a qubit drive of -3.0 dBm. Blue line: the qubit drive is detuned from the HBAR by $250$ kHz with the phonon mode in the non-lasing state, decaying on a time scale of the $\sim 200$ ns decay of the qubit. Red line: The qubit drive is tuned directly on resonance with the HBAR mode, exciting it into the lasing state, exhibiting a dramatically longer, non-exponential decay due to re-excitation from the coherently excited phonon mode. The master equation simulation (dashed black line) is plotted over the data.
Phonon Generation
Three distinct features can be observed within the two-tone spectrum when increasing the qubit drive power. First, the qubit linewidth is power-broadened Schuster et al. (2005); Gambetta et al. (2006). Large qubit drive powers increase the qubit decay rate due to increased stimulated emission. Less intuitive is the gradual disappearance and narrowing of the phonon-induced transparency window; see Fig. 2(c). The total linewidth of the transparency window is proportional to the phonon-qubit cooperativity and demonstrates an inverse dependence on the drive power; see Fig. 2(d). Such behavior can not be entirely explained by the power broadening of the qubit linewidth. At powers closer to the minimum of the curve Fig. 2(d), the narrowing of the transparency window linewidth can be associated with an amplification process triggered by the coupling to the driven qubit. Above such an upper threshold, the linewidth increases, a feature that can not be attributed to a suppression of the inverse Purcell effect due to the qubit broadening. These features can be understood by considering the schematic shown in Fig. 3(a).
The build-up of the phonon population can be understood by considering the different processes through which excitations can be transferred between the phonon and the qubit. At low driving powers, the linewidth broadening of the qubit due to the microwave drive is small, and the phonon mode and qubit are weakly interacting. The weak interaction combined with the microwave drive initiates the buildup of the phonon population. A balance between the rate at which the qubit is excited and de-excited, the rate at which excitations are transferred between qubit and phonon mode and the stochastic phonon losses yields a steady-state occupation of the phonon mode. For low drive powers and thus slow qubit excitation rates, phonon loss dominates, preventing the build-up of the phonon population. As the qubit excitation rate increases, transitions between states of the Jaynes-Cummings ladder with higher phonon numbers can be efficiently driven. Since the rate at which excitations are transferred between states of the Jaynes-Cummings ladder depends on the phonon number of the involved states, as the average population of the phonon mode increases, the rate at which excitations are transferred correspondingly increases; this is the origin of acoustic stimulated emission Hofheinz et al. (2008).
In the steady state, this process, balanced by stochastic phonon losses, results in a large phonon population. Such processes also provide an intuitive understanding of the reduced visibility of the transparency window. As the drive power increases, the linewidth of the qubit becomes increasingly broad, and the coupling between the qubit and phonon is not strong enough to overcome the qubit loss; therefore, the qubit spectroscopy provides less information about the phonon mode.
@@
At the highest qubit drive powers, the qubit linewidth has been increased such that the rapid decay of the qubit results in a quenching of phonon generation Mu and Savage (1992). This results in an upper threshold above which the phonon mode is no longer effectively excited. Above this threshold, the qubit undergoes rapid Rabi oscillations due to the strong microwave drive and, therefore, cannot exchange excitations with the phonon mode, reducing the phonon amplitude. The buildup of the phonon population, primarily due to stimulated emission from a single quantum emitter, including a reduced phonon mode linewidth and the upper threshold behavior, is similar to a single-atom laser. It should also be noted that the phonon mode statistics are no longer predicted to be that of a coherent state above the upper threshold. This process has been described previously in the context of single-atom photon lasers McKeever et al. (2003) and is captured by our theoretical description; see the supplementary text.
Figure 4: Unique signatures of predicted upper threshold. a Measured gated qubit ringdown for a qubit drive of -12.0, -8.0, -6.0, and -2.0 dBm. With increasing drive power, the gated qubit ringdowns increase in amplitude and duration as the phonon mode population increases. b Measured gated qubit ringdown for a qubit drive of -2.0, 2.0, 4.0, and 8.0 dBm. The qubit drive is tuned directly on resonance with the HBAR mode for all measurements. With increasing drive power, the gated qubit ringdowns decrease in amplitude and duration as the phonon mode population decreases above the self-quenching threshold. For all measurements in a and b, the qubit drive is tuned directly on resonance with the HBAR mode. c Simulated phonon state population $\langle\hat{b}^{\dagger}\hat{b}\rangle$ as a function of qubit drive power. The phonon population is at its maximum at a drive power of approximately -2.0 dBm and decreases for drive powers above the upper threshold. The colored data points indicate the corresponding trace colors in a and b.
@@
Directly probing the phonon mode is not possible in the current experimental configuration since the readout is performed via the two-level system and not through a propagating photon mode Kuang et al. (2023); Behrle et al. (2023). Direct measurements of the Rabi oscillations between the qubit and phonon state have been previously used to measure Fock-states in $\hbar$BAR devices Chu et al. (2018). However, this data would not be possible due to the short lifetime of our qubit and the multiplicity of simultaneous Rabi oscillation frequencies between two Fock states is given by $2g_{\rm N}=2\sqrt{N}g_{\rm qb}$, which scale with the phonon Fock number $N$.Chu et al. (2018) Instead, we rely on the mismatch between the decay rate of the phonon mode $\tau\sim 25$ $\mu$s and that of the qubit. Using gated two-tone ringdown measurements, we can distinguish pure qubit decay from qubit decay driven by a highly excited phonon state. If the phonon mode is highly excited—in the absence of an external drive—the coherent Jaynes-Cummings interaction will continually drive the qubit, resulting in an extended relaxation of the qubit population compared to its intrinsic relaxation rate.
@@
First, the drive power was set near the peak of the lasing amplitude and was detuned $250$ kHz above the HBAR resonance. The gated two-tone measurement was performed, and the blue data points in Fig. 3(b) show the resulting time domain measurement and a ringdown on the order of $\sim 200$ ns. This corresponds to the intrinsic $T_{1}$ decay of the transmon qubit. A second measurement was performed at the same drive power, but the drive was tuned resonant with the HBAR. A ringdown on the order of $25$ $\mu$s is observed for this configuration, represented by the red data points in Fig. 3(b). The extended ringdown confirms the highly excited phonon amplitude of the mechanical state. When the qubit drive is switched off, the phonon mode can exchange excitations with the qubit, continually re-exciting the qubit until the phonon mode has decayed back to its ground state. Moreover, near the peak phonon amplitude, in contrast to a thermal state, the coherent state generated by the lasing results in a qubit ringdown that is not exponential; instead, the qubit ringdown is approximately linear. This feature is captured by our numerical model, the dashed curve in Fig. 3(b). Moreover, our model also captures the ring-up of the qubit, which is described in the supplementary information.
@@
We can estimate the phonon population from our numerical model by fitting the spectroscopic and ringdown data. The estimated phonon population is shown in Fig. 4(c) as a function of qubit drive power. At a power of -2.0 dBm, the phonon population is nearly maximized, corresponding to the ringdown in Fig. 4(a,b), and the upper threshold is clearly visible as the phonon population rapidly reduces with increasing drive power. The upper threshold is experimentally confirmed by performing a set of gated ringdown measurements at a series of qubit drive powers. With increasing qubit drive power, the individual ringdown traces grow in amplitude and increase in duration, corresponding to the increasing phonon population, as shown in Fig. 4(a). At a drive power of -2.0 dBm, the phonon ringdown obtains its peak amplitude and duration, indicating a peak in the phonon population, shown in Fig. 4, which agrees with the minimum in the transparency window linewidth, shown in Fig. 2(b). Further increasing the qubit drive power beyond -2.0 dBm, the qubit ringdown decreases in amplitude and duration, a direct indication of the self-quenching, well described by our theoretical model and similar to what has been observed in single-atom lasers, see Fig. 4(b). The gated ringdown measurement demonstrates a clear upper threshold behavior and agrees with our numerical simulations and semi-classical analysis. We also notice that the qubit decay profile differs below and above the threshold for a given phonon population. Specifically, the decay is not exponential below the upper threshold, while well above the threshold, the decay is exponential. We associate such behavior with a change in the phonon state, which, according to our numerical simulations, is coherent below the upper threshold.
Discussion